# aoas/sm_enkf

### v2

@@ 3.2 Convergence of High Dimensional Ensemble Kalman Filter
Compared with the one-step results given in Theorem 3.1, the term related to the estimation error of $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the $h$-step analysis MSE is inflated with constants $(C_{t,h}^{A})^{2}$ which are related to $L_{t,h}$ and $A_{t,h}$ defined before (3.7). However, for any fixed $h$, $\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}\asymp\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}.$
@@ 4 Imperfect Models
Theorem 4.1 presents the one-step assimilation bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter when $\mathcal{M}_{t}$ is unknown. Denote
@@
$+\operatorname{tr}\big{(}n^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({\mathbf{H}}_{t}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}}){\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t})\big{)}$
$+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}-{\mathbf{H}}_{t}\mu_{\xi_{t}})\|^{2}+2\operatorname{tr}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}$ (4.1)
@@
Suppose Assumptions 1(i) or 1(ii), 2, 3, 4 hold and $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t-1}^{a}$ and covariance $\boldsymbol{\Sigma}_{t-1}^{a}$ at time $t-1$. Then, the forecast and analysis states of the EnKF satisfy
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=n^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}})+p^{-1}\|\mu_{\xi_{t}}\|^{2},$ (4.2)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=p^{-1}\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+V_{t}$ (4.3)
@@
where $\tilde{\Delta}_{\mathbf{K},t}=\mathbb{E}\|\tilde{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ and $\tilde{\mathbf{K}}_{t}$ is the estimate of $\mathbf{K}_{t}$ with true model $\mathcal{M}_{t}$.
@@
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}\text{ and }$ (4.4)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
@@
The above analysis shows that for both the fixed and diverging dimensions and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the assimilation MSE, which suggests that a good approximation of the model operator is a significant aspect of data assimilation. While this may be obvious, Theorem 4.1 describes exactly how it happens. The accumulation of the model approximation error is more pronounced in the $h-$step assimilation MSEs as shown in Theorem S.3.3 of the SM. Thus, an iterative HD-EnKF algorithm is introduced in S.1.2 to account for the model approximation error to some extent since most geophysical models are imperfect in practice, of which the main idea is to substitute $\hat{\mathbf{x}}_{t}^{f}$ in the original forecast sample covariance matrix by the analysis state $\hat{\mathbf{x}}_{t}^{a}$ in the iteration.
5 Simulation
@@
This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed HD-EnKF algorithms. These two models are often used for testing new approaches in atmospheric sciences for both data assimilation and the spatial-temporal forecasts. The proposed HD-EnKF methods are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models. Details of the proposed HD-EnKF algorithms are given in Section S.1 of the SM, including the consistent HD-EnKF update in S.1.1 and the iterative HD-EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators.
@@ 5.1 Lorenz-96 model
Lorenz-96 model (L96; Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere. It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation
${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),\,\text{ for }j=1,\ldots,p,$ (5.1)
@@
with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denotes the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $-\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is a forcing term.
@@
L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016). An approximation to the true state was simulated by (i) implementing the RK4 method with a finer time step, say $h=0.05$; and (ii) adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$. In our experiment, $\mathbf{w}_{t}\sim\mathrm{N}(0,\sigma_{0}\mathbf{I}_{p})$ and $\sigma_{0}$ were set as $0.1,0.25$ and $0.5$, respectively.
We experimented $p=40$ and $100$, under which the system behaves chaotically, since the L96 model behaves like chaos for $p\geq 7$ (Bedrossian et al., 2022). The true forcing was $F=8$. We also considered the misspecified model setting by making $F=6,7,9,10$ in the assimilation processes, corresponding to the imperfect models in Section 4. The initial state was $x_{1}^{(j)}=F$ for $j\neq 20$ and $x_{1}^{(20)}=F+0.001$, and the initial forecast ensemble was generated by $\mathbf{x}_{1,j}^{f}=\mathbf{x}_{1}+\mathrm{N}_{p}(0,0.1\mathbf{I}_{p})$ for $j=1,\ldots,n$.
@@
The observations $\mathbf{y}_{t}$ were made at $q$ randomly selected components of the state vector $\mathbf{x}_{t}$ with $q=30,40$ and $100$ while $p=40$ and $100$, respectively, followed by adding observation errors $\boldsymbol{\varepsilon}_{t}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{q}(0,\mathbf{R}_{t})$ to the true states. The $(i,j)-$element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,q-|i-j|\right\}}$, which prescribes spatially correlated observation errors. The observation operator $\mathbf{H}_{t}$ was a $q\times p$ matrix with the $(j,i)$ element being $1$ if and only if the $j-$th component of $\mathbf{y}_{t}$ was the observed $i-$th dimension of the state $\mathbf{x}_{t}$. We simulated observations every $4$ time steps for a total of $2000$ steps and reported the results for the last $1000$ steps to allow the system to settle down. The ensemble size $n$ was $20,30,60$ and $90$. All the experiments were repeated 500 times.
@@
We considered the HD-EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance to estimate the forecast error covariance and the inflation methods (Liang et al., 2012). The performances of these methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$. To gain accurate estimations of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance matrices to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman filter update formula to obtain $\mathbf{x}_{t}^{a}$.

### v3

@@ 3.2 Convergence of High Dimensional Ensemble Kalman Filter
Compared with the one-step results in Theorem 3.1, the term related to the estimation error in $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the $h$-step analysis MSE is inflated with constants $(C_{t,h}^{A})^{2}$ which are related to $L_{t,h}$ and $A_{t,h}$ defined after (3.6). This is because of the accumulation of the estimation error of $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the data assimilation, which can be seen from the proof of Theorem 3.2 in the SM. Using a similar analysis on the one-step results after Theorem 3.1 and according to (3.1), the $h$-step assimilation MSE between the analysis state of the EnKF $\hat{\mathbf{x}}_{t+h}^{a}$ and the $\mathbf{x}_{t+h}^{a}$ for $\boldsymbol{\Sigma}_{t}^{f}$ estimated by the four high dimensional estimators are, respectively,
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;(C_{t,h}^{A})^{2}{p}^{-1}q\big{(}{n}^{-1}{\log p}\big{)}^{\alpha/(\alpha+1)}\big{\}},$ (3.7)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;(C_{t,h}^{A})^{2}{p}^{-1}q(n^{-\frac{2\alpha}{2\alpha+1}}+{n}^{-1}{\log p})\big{\}},$ (3.8)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;(C_{t,h}^{A})^{2}{p}^{-1}qc_{0}^{2}(p)\big{(}n^{-1}{\log p}\big{)}^{(1-\gamma)}\big{\}},$ (3.9)
in which the estimation error of $\mathbf{K}_{k}$ are inflated by $(C_{t,h}^{A})^{2}$ compared with that in the upper bound of the one-step assimilation MSE given in (3.3) - (3.5) in Theorem 3.1. According to Remark 1, $A_{t,h}=\max_{t<k\leq t+h}{\left\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\right\|}<1$. If $L_{t,h}A_{t,h}<1$, then $C_{t,h}^{A}$ converges to a small constant and $L_{t,h}^{h}A_{t,h}^{h}\rightarrow 0$ as $h\rightarrow\infty$, although $C_{t,h}^{A}$ becomes larger along with the increase of $h$. Thus, $\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$. If $L_{t,h}A_{t,h}>1$, then $C_{t,h}^{A}$ increases a lot with the increase of $h$, which suggests the quality of the data assimilation deteriorates as $h$ increases, although $\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ for a fixed $h$. However, it is possible to make $L_{t,h}$ be smaller by shortening the forecast time step or the integration time step. This underscores the necessity of the high-quality or high dimensional estimators of the forecast error covariance, considering the accumulation of estimation errors in multiple-step data assimilation. The MSE between the forecast state of the EnKF $\hat{\mathbf{x}}_{t+h}^{f}$ and the Kalman Filter $\mathbf{x}_{t+h}^{f}$ for the banding, mid-banding, tapering and thresholding estimators of $\boldsymbol{\Sigma}_{t}^{f}$ are similar to (3.7)-(3.9) but with slightly different inflation constants.
@@ 4 Imperfect Models
Theorem 4.1 gives the one-step assimilation bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter when $\mathcal{M}_{t}$ is unknown. Denote
@@
$+\operatorname{tr}\big{(}n^{-1}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})^{\mathrm{T}}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})({\mathbf{H}}_{t}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}}){\mathbf{H}}_{t}^{\mathrm{T}}+\mathbf{R}_{t})\big{)}$ (4.1)
$+\mathbb{E}\|(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})(\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}-{\mathbf{H}}_{t}\mu_{\xi_{t}})\|^{2}+2\operatorname{tr}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}$
@@
[…]
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f}\|^{2}|\mathbf{y}_{1:t-1}\}=n^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f}+\Xi_{\xi_{t}})+p^{-1}\|\mu_{\xi_{t}}\|^{2},$ (4.2)
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=p^{-1}\operatorname{tr}\big{(}n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+V_{t}$ (4.3)
@@
where $\tilde{\Delta}_{\mathbf{K},t}=\mathbb{E}\|\tilde{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ and $\tilde{\mathbf{K}}_{t}$ is the estimated $\mathbf{K}_{t}$ with true model $\mathcal{M}_{t}$.
@@
$\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}$ (4.4)
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
@@
$\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O\big{(}n^{-1}+\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}(\|\Xi_{\xi_{t}}\|^{2}+n^{-1})\big{)}=O\big{(}n^{-1}+\|\Xi_{\xi_{t}}\|^{2})\big{)}.$ (4.5)
Thus, if $\|\Xi_{\xi_{t}}\|^{2}=O(n^{-1})$, then $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O(n^{-1})$, that is, the model approximation error does not alter the order of magnitude of the assimilation MSEs of the EnKF. However, if $n\|\Xi_{\xi_{t}}\|^{2}\rightarrow\infty$, then $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O(\|\Xi_{\xi_{t}}\|^{2})$, which means the model approximation error dominates the assimilation MSE. When $p,q\rightarrow\infty$, $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}\asymp\max\{n^{-1},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
[…]
$\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;\|\Xi_{\xi_{t}}\|^{2}\}\text{ and }$ (4.6)
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;p^{-1}\|(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\mu_{\xi_{t}}\|^{2},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$
@@
The above analysis shows that for both the fixed and diverging dimensions and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the assimilation MSE, which suggests that a good approximation of the model operator is a significant aspect of data assimilation. While this may be obvious, Theorem 4.1 describes exactly how it happens.
The accumulation of the model approximation error is more pronounced in the $h$-step assimilation MSEs as shown in Theorem 4.2 and more details of the $h$-step assimilation in the presence of the model error can be found in Theorem S.3.4 of the SM. Define $D_{t,h}=\max_{t<k\leq t+h}{\left\|\mathbf{I}-\check{\mathbf{K}}_{k}\mathbf{H}_{k}\right\|}$, where ${\check{\mathbf{K}}}_{k}=(\boldsymbol{\Sigma}_{k}^{f}+\Xi_{k})\mathbf{H}_{k}^{\mathrm{T}}\big{(}\mathbf{H}_{k}(\boldsymbol{\Sigma}_{k}^{f}+\Xi_{\xi_{k}})\mathbf{H}_{k}^{\mathrm{T}}+\mathbf{R}_{k}\big{)}^{-1}$, which is the underlying Kalman Gain matrix under the imperfect model. Denote a non-stochastic constant $C_{t,h}^{D}=(1-L_{t,h}^{h}D_{t,h}^{h})/(1-L_{t,h}D_{t,h})$.
Theorem 4.2.
[…]
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f}\|^{2}|\mathbf{y}_{1:t+h-1}\}$
$=O\Big{(}n^{-1}+p^{-1}\|\mu_{\xi_{t+h}}\|^{2}+p^{-1}L_{t,h}^{2}(C_{t,h-1}^{D})^{2}\max_{t<k\leq t+h-1}\big{\{}\Delta_{\check{\mathbf{K}}_{k}}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}-\mathbf{H}_{k}\mu_{\xi_{k}}\|^{2}\big{\}}\Big{)},$
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}$
$=O\Big{(}n^{-1}+p^{-1}\|(\mathbf{I}-\mathbf{K}_{t+h}\mathbf{H}_{t+h})\mu_{\xi_{t+h}}\|^{2}+p^{-1}\|\Xi_{\xi_{t+h}}\|^{2}\|\mathbf{y}_{t+h}-\mathbf{H}_{t+h}\mathbf{x}_{t+h}^{f}\|^{2}$
$\quad+p^{-1}(C_{t,h}^{D})^{2}\max_{t<k\leq t+h}\big{\{}\Delta_{\check{\mathbf{K}}_{k}}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}-\mathbf{H}_{k}\mu_{\xi_{k}}\|^{2}\big{\}}\Big{)}.$
Compared with the result in Theorem 3.2 under the correct model, there are extra terms related to $D_{t,h}=O(\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|+\|\Xi_{\xi_{k}}\|\|\mathbf{H}_{k}\|)$, $\|\mu_{\xi_{t+h}}\|$ and $\|\Xi_{\xi_{t+h}}\|$ resulting from the accumulation of the approximation error of the model operator $\mathcal{M}_{t}$ for the $h$-step assimilation MSE in Theorem 4.2. Compared with the one-step assimilation MSE in Theorem 4.1 under the imperfect models, the error term related to the approximation error of $\mathcal{M}_{t}$ is inflated by $(C_{t,h}^{D})^{2}$. The $h$-step assimilation MSE for the imperfect models under the fixed and high dimensions are, respectively,
$\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\max\big{\{}\big{(}1+(C_{t,h}^{D})^{2}\big{)}n^{-1},\;\|\mu_{\xi_{t+h}}\|^{2},\;\|\Xi_{\xi_{t+h}}\|^{2}\big{\}},$
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a}\|^{2}|\mathbf{y}_{1:t+h}\}\asymp\max\big{\{}n^{-1},\;p^{-1}\|\mu_{\xi_{t+h}}\|^{2},\;p^{-1}{q}\|\Xi_{\xi_{t+h}}\|^{2},\;(C_{t,h}^{D})^{2}p^{-1}{q}\Delta_{\check{\mathbf{K}}_{k}}\big{\}}.$ (4.7)
Specifically, for the high dimensional case, the MSEs of the EnKF analysis state $\hat{\mathbf{x}}_{t+h}^{a}$ to the Kalman Filter $\mathbf{x}_{t+h}^{a}$ for the $h$-step assimilation using the banding or mid-banding, tapering and thresholding estimates of $\boldsymbol{\Sigma}_{t}^{f}$ can be obtained by replacing $\Delta_{\check{\mathbf{K}}_{k}}$ in equation (4.7) with the specific convergence rates $({n}^{-1}{\log p})^{\alpha/(\alpha+1)}$, $n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p}$ and $c_{0}^{2}(p)(n^{-1}{\log p})^{(1-\gamma)}$, respectively, according to (3.1). For the $h$-step assimilation, the approximation error of the model operator $\mathcal{M}_{t}$ has a stronger influence on the MSE of the data assimilation of the EnKF resulting from the accumulation of the approximation error. This highlights that a good approximation of the model operator is significant to data assimilation.
In summary, the approximation error associated with the model operator $\mathcal{M}_{t}$ significantly affects the MSE in the assimilation process of the EnKF, particularly in scenarios involving multiple-step ahead data assimilation. Thus, an iterative HD-EnKF algorithm is introduced in S.1.2 of the SM to account for the model approximation error to some extent since most geophysical models are imperfect in practice, of which the main idea is to substitute $\hat{\mathbf{x}}_{t}^{f}$ in the original forecast sample covariance matrix by the analysis state $\hat{\mathbf{x}}_{t}^{a}$ in the iteration.
5 Simulation Studies
@@
This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed HD-EnKF algorithms. These two models are often used for testing new approaches in atmospheric sciences for data assimilation and the spatial-temporal forecasts. The proposed HD-EnKFs are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models. Details of the proposed HD-EnKF algorithms are given in Section S.1 of the SM, including the consistent HD-EnKF update in S.1.1 and the iterative HD-EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators. The iterative HD-EnKF algorithm in S.1.2 of the SM is to account for the model approximation error in the imperfect models.
@@ 5.1 Lorenz-96 model
Lorenz-96 model (L96; Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere. It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation, for $j=1,\ldots,p$,
${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),\,$ (5.1)
@@
with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denotes the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is a forcing term.
@@
L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016), which approximates the model as $\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)=\mathbf{x}_{t}+h(k_{1}+2k_{2}+2k_{3}+k_{4})/{6},$ where $k_{1}=\mathbf{f}\left(\mathbf{x}_{t},t\right)$, $k_{2}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{1}/2,t+h/2\right)$, $k_{3}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{2}/2,t+h/2\right)$, $k_{4}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{3},t+h\right)$ and $\mathbf{f}=\left(f_{1},f_{2},\dots,f_{p}\right)^{\mathrm{T}}$. An approximation to the true state was simulated by (i) implementing the RK4 method with a finer time step, say $h=0.05$; and (ii) adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$. In our experiment, $\mathbf{w}_{t}\sim\mathrm{N}(0,\sigma_{0}\mathbf{I}_{p})$ and $\sigma_{0}$ were set as $0.1,0.25$ and $0.5$, respectively. We experimented $p=40$ and $100$, under which the system behaves chaotically, since the L96 model behaves like chaos for $p\geq 7$ (Bedrossian et al., 2022). The true forcing was $F=8$. We also considered the misspecified model setting by making $F=6,7,9,10$ in the assimilation processes, corresponding to the imperfect models in Section 4. The initial state was $x_{1}^{(j)}=F$ for $j\neq 20$ and $x_{1}^{(20)}=F+0.001$, and the initial forecast ensemble was generated by $\mathbf{x}_{1,j}^{f}=\mathbf{x}_{1}+\mathrm{N}_{p}(0,0.1\mathbf{I}_{p})$ for $j=1,\ldots,n$.
@@
The observations $\mathbf{y}_{t}$ were made at $q$ randomly selected components of the state vector $\mathbf{x}_{t}$ with $q=30,40$ and $100$ while $p=40$ and $100$, respectively, followed by adding observation errors $\boldsymbol{\varepsilon}_{t}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{q}(0,\mathbf{R}_{t})$ to the true states. The $(i,j)$-element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,q-|i-j|\right\}}$, which prescribes spatially correlated observation errors. The observation operator $\mathbf{H}_{t}$ was a $q\times p$ matrix with the $(j,i)$ element being $1$ if the $j$-th component of $\mathbf{y}_{t}$ was the observed $i$-th dimension of the state $\mathbf{x}_{t}$. We simulated observations every $4$ time steps for a total of $2000$ steps and reported the results for the last $1000$ steps to allow the system to settle down. The ensemble size $n$ was $20,30,60$ and $90$. All the experiments were repeated 500 times.
@@
We considered the HD-EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance and the inflation methods (Liang et al., 2012). The performances of these methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$. To gain accurate estimations of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman Filter update formula to obtain $\mathbf{x}_{t}^{a}$.